Systems Staff Meeting Notes 2/14/00

Systems Staff Meeting Notes
February 14, 2000

Present: Rich Bennett, Will Chaney, Bill Covey, Debra Harris, Winston Harris, Bobbie Meng, Everett Phillips, Suzy Shaw, Hilton Soto, Ying Tang
Announcements
  1. West feeder switch just died, it was down about five minutes.
  2. Bill's work station crashed on Friday, luckily he had all his files backed up on the server.
  3. The last LMG meeting was unusually rich - many prophesies on the table.
  4. Ricky the Rat update - he nibbled on network wires in the server room recently. Glue traps that were placed in the conference room have been removed.
  5. Had a termite swarm in Systems last week.
Review Minutes of 1/31/00
  1. DOS attacks - following our discussion last week, the whole world became aware of the problem. University of California campuses were used as points of origin for attacks. UF still has no solution to the problem. Washington has scheduled a big conference on the DOS attacks - not much hope for a real solution.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Eight Cows received - should have come with SCSI read/write CD ROM drives but showed up with IDE. Have not been able to contact Gateway - we hope to switch out the drives.
LMG Recap
  1. Dale thinks we're close to starting the new building in the back of West - 5 stories high, 20,000 square feet - compact storing throughout. We have the chance to tap 7-10 million dollars in PECO funds. Total cost could be about 30 million dollars. The proposal will be up for approval in August, construction could possibly start in June 2001.
  2. Rat button - Rob Roberts & Cathy Mook have devised a pest tracking database that will be put on the server and then optionally downloaded to individual desktops. They are trying to improve the speed & efficiency in reporting vermin sightings. The system will enable prompt email notification to pest control, recording the person who sent it so that direct contact can be made for more details.
  3. Cathy gave a presentation on shelving non-paper materials - computer readable items. They will take gizmos (like anti-pilfering devices for CDs you find in stores) and shelve them as they do books.
  4. Trudi discussed how Library Personnel has decided to let veteran staff members attend orientation sessions. There was much banter about the lunch portion of the orientation - the budget is really too limited to provide it for all. The compromise was to have two orientation sessions - one where the newcomers get lunch and another where the old timers do not.
  5. Web Police - a committee will be scanning our websites to make sure they correspond to the UF standards of conformity. Official templates for UF in FrontPage will be made available.
  6. FCLA - they are holding a contest to come up with an umbrella name for LUIS & the whole digital system. The prize being offered is a laptop. Multiple entrees are allowed. Contact Carol Turner for further information.
  7. MSL reports that the lighting on the first floor should be fixed next week.
  8. The Foundation has put up web pages to solicit contributions and the Library has received $100,000 in the first donation.
Networking
  1. Will reports that one of the switches in West was not working this morning. It might be related to the power fluctuations we had in the building last week. It went down at 11 AM, but was back up by 11:15 AM.
  2. There was a general campus outage at approximately the same time this morning.
  3. Bill & Will have a meeting with NERDC scheduled for Wednesday morning regarding networking.
  4. No signs yet of wiring crews in West. Should be working on new com closets, phone wiring and a few drops. Should be completed by March.
  5. NT9 was copying files again - appears to be working okay. It lost a RAID drive, Will spent time rebuilding.
  6. Proxy server overloads in the morning - predominantly from MSL - the problem clears itself up after a few minutes. The problem could be the switches in MSL, the 1900's need to be replaced. The 1900's are reportedly not performing up to spec. Replacing them in Resource Services took care of the timeout problems. The whole notion of a proxy server is that it should be able to take a zillion hits at once. We'll try swapping out the switches.
Software Projects Status
  1. Ying & Hilton have been working on the filtering of web pages to prevent hacking attempts into our system. Hilton came up with a procedure, although we are still looking to be assured of the best solution.
  2. Ying is finishing up OPS statistic display, as well as some repairs on IR & wiring.
Calendar
  1. Today is Valentines Day.
  2. Winston out Feb. 16-18.
  3. Hilton out Feb. 17.


Back to the Systems Homepage

Last updated February 21, 2000
by Debra Harris
debharr@mail.uflib.ufl.edu